### v1

# Correctable Landmark Discovery via Large Models for Vision-Language Navigation

Abstract
Vision-Language Navigation (VLN) requires the agent to follow language instructions to reach a target position. A key factor for successful navigation is to align the landmarks implied in the instruction with diverse visual observations. However, previous VLN agents fail to perform accurate modality alignment especially in unexplored scenes, since they learn from limited navigation data and lack sufficient open-world alignment knowledge. In this work, we propose a new VLN paradigm, called COrrectable LaNdmark DiScOvery via Large ModEls (CONSOLE). In CONSOLE, we cast VLN as an open-world sequential landmark discovery problem, by introducing a novel correctable landmark discovery scheme based on two large models ChatGPT and CLIP. Specifically, we use ChatGPT to provide rich open-world landmark cooccurrence commonsense, and conduct CLIP-driven landmark discovery based on these commonsense priors. To mitigate the noise in the priors due to the lack of visual constraints, we introduce a learnable cooccurrence scoring module, which corrects the importance of each cooccurrence according to actual observations for accurate landmark discovery. We further design an observation enhancement strategy for an elegant combination of our framework with different VLN agents, where we utilize the corrected landmark features to obtain enhanced observation features for action decision. Extensive experimental results on multiple popular VLN benchmarks (R2R, REVERIE, R4R, RxR) show the significant superiority of CONSOLE over strong baselines. Especially, our CONSOLE establishes the new state-of-the-art results on R2R and R4R in unseen scenarios. Code is available at https://github.com/expectorlin/CONSOLE.
Index Terms:
Vision-language navigation, open-world landmark discovery, large language models
A Introduction
Vision-and-Language Navigation (VLN) [1, 2, 3, 4, 5], one of the most representative embodied AI tasks, requires an agent to navigate through complicated visual environments to a goal position following a given natural language instruction. It has led to a wide range of research recently since an instruction-following navigation agent is more flexible and practical in real-world applications. For successful navigation, the agent needs to accurately understand the instruction intention and align the landmarks implied in the instruction to sequential visual observations.
Early VLN methods utilize different data augmentation strategies [6, 7, 8], efficient learning paradigms [9, 10, 11], and dedicated model architectures [12, 13, 14] to learn useful modality alignment knowledge from limited human annotated navigation data. With the development of cross-modal pretraining models, increasing pretraining-based VLN approaches are developed [15, 16] and have shown great progress in improving the modality alignment ability of VLN agents. However, due to the limited scale and diversity of pretraining and navigational data, these approaches still cannot generalize well to unseen navigation scenarios, which usually require rich open-world alignment knowledge, e.g., recognizing the unseen landmark, to realize accurate instruction following.
Figure 1: Correctable landmark discovery of CONSOLE. We provide customized prompts for ChatGPT to generate landmark cooccurrence priors. Then we introduce a learnable cooccurrence scoring module to conduct CLIP-driven correctable landmark discovery based on the priors. The landmark/cooccurrence with bigger font size has a higher score.
Recent studies have shown that employing large models, such as Large Language Models (LLMs, e.g., GPT-3 [17]) and large Vision-Language Models (VLMs, e.g., CLIP [18]), for embodied AI tasks is a very promising way to improve the task completion [19, 20, 21], since the large models are pretrained on ultra-large-scale web corpora and capture vast amount of open-world knowledge which is helpful for embodied AI tasks. However, constrained by the large domain gap between those training corpora and embodied AI task datasets, a direct introduction of large models for embodied AI tasks may bring unexpected noise [19, 22]. Therefore, how to effectively activate and more importantly, utilize the task-related knowledge from large models to assist embodied agents, has been paid more and more attention in the embodied AI area.
In this paper, we propose a new paradigm called COrrectable LaNdmark DiScOvery via Large ModEls (CONSOLE), where we activate and utilize rich open-world alignment knowledge stored in two large models ChatGPT [23] and CLIP [18] for assisting the VLN task in a correctable way. Through CONSOLE, we cast VLN as an open-world sequential landmark discovery problem for facilitating action decision. Our approach is inspired by the navigation commonsense, that is, the landmark cooccurrence is helpful for locating a target landmark. For example, a fridge is likely to be placed near a stove. When a mentioned landmark in the instruction is hardly observable in the views, locating the landmark cooccurrence (which is not mentioned in the instruction) is an effective alternative of locating the mentioned landmark to make correct navigation decisions. Being trained using ultra-large-scale corpora, ChatGPT captures sufficient open-world landmark cooccurrence commonsense. However, due to the lack of visual constraints, these commonsense priors may be inconsistent with actual navigation scenes and mislead the agent to making wrong action decision. To mitigate the impact of the noisy priors, we introduce a CLIP-driven correctable landmark discovery scheme, where the priors are corrected in a learnable way according to actual observations. A consistency loss is introduced to constrain the prior correction using the navigation supervision. As a result, at each navigation timestep, the importance of each cooccurrence is re-ranked based on the current observation to suppress the noisy cooccurrences for accurate navigation decision. As shown in Fig. 1, through correctable landmark discovery, CONSOLE successfully highlights the important landmark and cooccurrence (hall and artwork) for making correct action decision while getting rid of the impact of the noisy ones (rug and chandelier).
Our approach contains three core components: 1) Landmark cooccurrence prior generation: We design customized prompts for ChatGPT to extract landmark cooccurrence priors, including navigational sequential landmarks and open-world landmark cooccurrences. 2) Correctable landmark discovery: We introduce a learnable cooccurrence scoring module constrained by a consistency loss for accurate landmark discovery based on the priors. The scoring module ranks the importance of each cooccurrence according to dynamic navigation observations. 3) Observation enhancement: We use corrected landmark features for enhancing observation features to combine the landmark discovery framework with different VLN agents. The enhanced observation features are adopted for final action decision.
We evaluate CONSOLE on multiple mainstream VLN benchmarks, including R2R [1], REVERIE [2], R4R [3], and RxR [4]. The experimental results show that CONSOLE outperforms strong baselines [16, 24] on all benchmarks. Moreover, CONSOLE achieves the new state-of-the-art on R2R and R4R in unseen scenarios. We conduct extensive insightful visualization to verify the effectiveness and necessity of different components in CONSOLE, such as the learnable scoring module and the customized designs for the prompts. The visualization comprehensively reveals how CONSOLE successfully activates and utilizes the helpful navigation knowledge in large models for assisting action decision. We also present both quantitative and qualitative comparison of different LLMs to deeply analyze their potential in assisting VLN tasks. We believe that our work will provide a non-trivial reference in introducing large models for assisting embodied AI tasks in a noise-suppressed way.
To summarize, our main contributions are:
  • We propose a novel VLN paradigm called CONSOLE, where we activate and utilize the rich open-world knowledge in large models in a correctable way, which effectively assists VLN while mitigating the misleading brought by large models.
  • We cast VLN as an open-world sequential landmark discovery problem to effectively harvest the knowledge helpful for VLN from both LLM and VLM. Built upon an observation enhancement strategy, our correctable landmark discovery scheme can be combined with different VLN models elegantly.
  • CONSOLE outperforms strong baselines on multiple popular VLN benchmarks, including R2R, REVERIE, R4R and RxR. It also establishes new SOTA results on R2R and R4R in unseen scenarios.
B Related Work
B.1 Vision-Language Navigation.
In Vision-Language Navigation (VLN), the agent needs to move across complicated visual environments to reach a goal location following human instructions [1, 4, 2]. Therefore, accurate modality alignment is crucial for successful navigation. Early methods mainly design diverse data augmentation techniques [6, 7, 8, 25], efficient learning mechanisms [26, 9, 11, 10, 27, 28, 29], and useful model architectures [26, 12, 14, 30] for learning useful modality alignment knowledge. EnvDropout [6] creates the augmentation data by mimicking unseen scenes via an environmental dropout scheme. AuxRN [9] designs multiple self-supervised auxiliary training tasks to exploit rich semantic information in the environment. OAAM [30] designs a new model structure and loss for separately processing the mentioned objects and actions in the instructions.
Recent VLN works employ cross-modal pretraining paradigms and models for enhancing the modality alignment [15, 16, 31, 24, 32, 33, 34]. HAMT [16] introduces a history-aware multimodal transformer to encode the long-horizon navigation history. DUET [24] builds a dual-scale graph transformer to jointly perform long-term action planning and fine-grained cross-modal understanding. HOP [33] introduces a new history-and-order aware pretraining paradigm for encouraging the learning of spatio-temporal multimodal correspondence.
In this paper, we resort to existing powerful large models to cast VLN as an open-world sequential landmark discovery problem, for facilitating action decision in unseen scenes. we introduce a novel correctable landmark discovery scheme to correct the priors provided by large models according to actual observations in a learnable way. By using corrected landmark features for enhancing the observation features, we fulfill an elegant combination of our framework with different VLN agents.
Although some VLN approaches have also explored to decouple the landmark discovery for better cross-modality alignment [30, 35, 31], they tend to use the limited information in the instruction for aligning with the visual observation. In contrast, our method uses large models to provide rich commonsense prior accompanied by a powerful learnable prior correction paradigm, which enables the agent to learn much more open-world alignment knowledge for successful navigation.
B.2 LLMs & VLMs for Embodied Tasks.
Recent studies have shown that LLMs and VLMs can connect human instructions to robot actions and facilitate the alignment between instructions and scenes [36, 19, 20, 37, 38, 39, 22, 40, 41, 42], thanks to rich real-world and open-world knowledge stored in them. Planner [36] employs an LLM to decompose the abstract task instructions to concrete step-by-step executable actions for embodied agents. SayCan [19] further uses an LLM for assisting the action decision by grounding the LLM through value functions of pretrained robotic skills. Inner Monologue [20] introduces a VLM to provide environment feedbacks for an LLM to enable it to recover from failure in robotic control scenes. LM-Nav [37] combines an LLM, a VLM, and a visual navigation model to enable long-horizon instruction following without requiring user-annotated navigational data. However, directly using an LLM and/or a VLM for assisting VLN may introduce unexpected noise that confuses the action decision, due to the limited action space of VLN and the large domain gap between the training corpora of large models and the VLN dataset.
In this work, we activate open-world alignment knowledge from both LLM and VLM to realize an open-world landmark discovery paradigm for assisting VLN. To the best of our knowledge, we are the first to harvest landmark cooccurrence knowledge from LLM to assist the VLN task. Through a correctable landmark discovery scheme, the noise in the priors can be effectively suppressed for accurate action decision.
A concurrent work ESC [43] also introduces the landmark coocurrence knowledge from LLMs for assisting the zero-shot object navigation task. However, our CONSOLE is different from it in the following two key aspects: 1) we introduce a learnable cooccurrence scoring module to effectively mitigate the noise in the knowledge, which does not require complicated exploration procedure like Zhou et al. [43] and hardly impacts the inference speed. This module can also be easily optimized through navigation supervision. 2) Different from Zhou et al. [43] that focus on the zero-shot setting, our work proposes a trainable scheme to elegantly combine LLM, VLM, and the VLN agent, which achieves significantly higher navigation performance than the zero-shot setting.
C Preliminaries
C.1 Problem Setup
In the discrete VLN task, an agent is initialized at a starting node and needs to explore the navigation connectivity graph $\mathcal{G}=(V,E)$ to reach a target node, following a language instruction $I$. $V$ and $E$ represent the nodes and edges in the navigation connectivity graph. At each timestep $t$, the agent receives a panoramic observation $O_{t}$ of its current node. $O_{t}$ composed of $N_{o}$ single-view observations $O_{t,n}$, i.e., $O_{t}=\{O_{t,n}\}_{n=1}^{N_{o}}$. Each $O_{t,n}$ contains an RGB image $B_{t,n}$ and the directional information. With the instructions $I$ and current visual observations $O_{t}$, the agent infers the action $\mathbf{a}_{t}$ for each step $t$ from the candidate navigable views $\mathcal{N}(O_{t})\in O_{t}$, which corresponds to the neighbor nodes of the current node.
In this paper, we propose to cast VLN as an open-world landmark discovery problem to introduce large models for assisting VLN tasks. We define the problem in the following. Specifically, the goal of the open-world sequential landmark discovery is to discover the mentioned landmarks that the agent may not see before sequentially to complete the navigation. Denote the landmark list extracted from the instruction $I$ as $U^{la}$. At each timestep $t$, the agent first needs to predict the landmark $U^{la}_{t}$ which is needed to find currently from the landmark list $U^{la}$. Then, it discovers $U^{la}_{t}$ in the set of navigable viewpoints and makes the action prediction $\mathbf{a}_{t}$ by selecting the navigable viewpoint containing $U^{la}_{t}$.
C.2 Large Models
Benefiting from ultra-large-scale web corpora, large pretrained models have shown excellent generalization abilities to various tasks due to their vast knowledge storage. In CONSOLE, we resort to two powerful large models ChatGPT [23] and CLIP [18], to provide rich open-world alignment knowledge for facilitating action decision.
ChatGPT [23] is a powerful text generation model that shows great in-context reasoning ability, i.e., given with customized prompts with a few task examples, its internalized knowledge about some specific task can be activated through text completion.
CLIP [18] is a vision-language model pretrained on 400M image-text pairs collected from the web. It utilizes a ViT [44] or a ResNet-50 [45] as the image encoder and a Transformer [46] as the text encoder. By conducting a simple similarity calculation between images and textual class descriptions, it shows powerful zero-shot object recognition ability in diverse datasets [18].
C.3 Baseline VLN Agents
We choose two strong baselines HAMT [16] and DUET [24] to verify the effectiveness of CONSOLE. In this section, we briefly describe one baseline HAMT. In HAMT, the agent receives the instruction $I$, the observation $O_{t}$, and the navigation history $H_{t}$ at each timestep $t$. For simplicity, we omit the encoding process for the navigation history $H_{t}$. Specifically, a text encoder $E^{I}(\cdot)$ and a vision encoder $E^{V}(\cdot)$ are employed to obtain the instruction feature $\mathbf{f}_{I}$ and the observation feature $\mathbf{f}_{O_{t}}$, respectively:
$\mathbf{f}_{I}=E^{I}(I),\quad\mathbf{f}_{O_{t}}=E^{V}(O_{t}).$ (1)
Then $\mathbf{f}_{I}$ and $\mathbf{f}_{O_{t}}$ are updated through a cross-modal Transformer encoder $E^{c}(\cdot)$:
$\mathbf{\tilde{f}}_{I}^{t},\mathbf{\tilde{f}}_{O_{t}}=E^{c}(\mathbf{f}_{I},\mathbf{f}_{O_{t}}).$ (2)
The action prediction probability $\mathbf{a}_{t}$ is generated through an action prediction module $E^{a}(\cdot)$ based on $\mathbf{\tilde{f}}_{I}^{t}$ and $\mathbf{\tilde{f}}_{O_{t}}$:
$\mathbf{a}_{t}=E^{a}(\mathbf{\tilde{f}}_{I}^{t},\mathbf{\tilde{f}}_{O_{t}}).$ (3)
The action decision is optimized by the navigation loss $\mathcal{L}_{\mathrm{nav}}$, which contains an imitation learning loss $\mathcal{L}_{\mathrm{IL}}$ and a reinforcement learning loss $\mathcal{L}_{\mathrm{RL}}$ [16] :
$\mathcal{L}_{\mathrm{IL}}=\sum_{t}-\mathbf{a}_{t}^{*}\mathrm{log}(\mathbf{a}_{t}),$ (4)
$\mathcal{L}_{\mathrm{RL}}=\sum_{t}-\mathbf{a}_{t}^{s}\mathrm{log}(\mathbf{a}_{t})A_{t},$ (5)
$\mathcal{L}_{\mathrm{nav}}=\mathcal{L}_{\mathrm{IL}}+\lambda\mathcal{L}_{\mathrm{RL}},$ (6)
where $\mathbf{a}_{t}^{*}$ is the teacher action of the ground-truth path at timestep $t$, $\mathbf{a}_{t}^{s}$ is the sampled action from the agent action prediction $\mathbf{a}_{t}$, $A_{t}$ is the advantage calculated by A2C algorithm [47], and $\lambda$ is a balance factor.
In CONSOLE, we use the corrected landmark features $\mathbf{f}_{U_{t}}$ and the observation features $\mathbf{f}_{O_{t}}$ to obtain the enhanced observation features $\mathbf{f^{\prime}}_{O_{t}}$ at each timestep $t$. Then $\mathbf{f^{\prime}}_{O_{t}}$ are fed into $E^{c}(\cdot)$ and $E^{a}(\cdot)$ to get final action prediction $\mathbf{\tilde{a}}_{t}$.
TABLE I: Notation summarization of variables.
Notations Variables
$I$ instruction
$O_{t}$ observation
$U$ landmark cooccurrence prior
$z$ shifting pointer
$U^{la}_{t}$ current important landmark
$s^{la}$ landmark score
$\{s_{i}^{co}\}_{i=1}^{N^{co}}$ cooccurrence score
$\mathbf{f}_{I}$ instruction feature
$\mathbf{f}_{O_{t}}$ observation feature
$\mathbf{f^{\prime}}_{O_{t}}$ enhanced observation feature
$\mathbf{f}_{S_{t}}$ state feature
$\mathbf{f}_{U_{t}}$ corrected landmark feature
$\mathcal{L}_{\mathrm{nav}}$ navigation loss
$\mathcal{L}_{\mathrm{cs}}$ consistency loss
$\mathcal{L}_{\mathrm{ct}}$ contrastive loss
Figure 2: Overview of CONSOLE. Before navigation, the landmark cooccurrence priors $U$ are obtained through the landmark cooccurrence prior generation module. At navigation timestep $t$, the agent conducts the correctable landmark discovery based on the landmark shifting, the landmark discovery, and the learnable cooccurrence scoring. An observation enhancing module is introduced to enhance the observation features $\mathbf{f}_{O_{t}}$ using the corrected landmark features $\mathbf{f}_{U_{t}}$. And the enhanced observations $\mathbf{f^{\prime}}_{O_{t}}$ are used for action decision. Besides the navigation loss $\mathcal{L}_{\mathrm{nav}}$, we introduce the consistency loss $\mathcal{L}_{\mathrm{cs}}$ and the contrastive loss $\mathcal{L}_{\mathrm{ct}}$ for optimization.
D Method
The overview of CONSOLE is presented in Fig. 2. For a given instruction $I$, we first conduct the landmark cooccurrence prior generation to obtain the landmark cooccurrence priors $U$ (Sec. D.1). Then, at each timestep $t$, we perform the correctable landmark discovery through the landmark shifting, the landmark discovery, and the learnable cooccurrence scoring, based on the priors $U$, the instruction $I$, and the observations $O_{t}$ (Sec. D.2). After the correctable landmark discovery, we obtain the corrected landmark feature $\mathbf{f}_{U_{t}}$and use it to enhance the observation feature $\mathbf{f}_{O_{t}}$ (Sec. D.3). We adopt the enhanced observation feature $\mathbf{f^{\prime}}_{O_{t}}$ for predicting the final action $\mathbf{\tilde{a}}_{t}$. The total training objective of CONSOLE contains the navigation loss $\mathcal{L}_{\mathrm{nav}}$ and two additional introduced losses, i.e., a consistency loss $\mathcal{L}_{\mathrm{cs}}$ and a contrastive loss $\mathcal{L}_{\mathrm{ct}}$ (Sec. D.4). For facilitating reading, we list the crucial variables in CONSOLE in Table I.
D.1 Landmark Cooccurrence Prior Generation
The landmark cooccurrence prior contains the following two elements: 1) Navigational Sequential Landmarks: Due to the complexity of natural language, the order of landmarks appearing in the instruction may be sometimes inconsistent with the intended order to be searched during navigation. Therefore, extracting navigational sequential landmarks is crucial for the landmark discovery. 2) Landmark Cooccurrence: the open-world landmark cooccurrence knowledge, e.g., a counter may appear in a kitchen island (Fig. 2), can greatly help the alignment. We design customized prompts with heuristic task examples to obtain these two kinds of priors.
Navigational Sequential Landmark Extraction. An ideal extraction of landmarks requires: (1) extraction of all mentioned landmarks; (2) successful extraction of landmarks with complex descriptions (e.g., landmarks with adjectives or phrases); (3) avoidance of abstract nouns (e.g., right); (4) navigational order instead of textual order. To ensure the coverage for diverse descriptions of landmarks for accurate extraction, we pre-define 5 different task examples for ChatGPT. Taking REVERIE as an example, one task example is as follows:
Instruction: Go to the lounge on the first level and bring the trinket that’s sitting on the fireplace.
Landmarks:
1.first level;
2.lounge;
3.fireplace;
4.trinket.
Through such example, we can let ChatGPT know that it should extract navigational sequential landmarks rather than textual sequential landmarks (1.lounge; 2.first level; 3.trinket; 4.fireplace). Moreover, it should extract “first level” as the landmark rather than “level”. After in-context learning with task examples, we can ask ChatGPT to extract the landmarks of a specific instruction $I$ with the following prompt:
Instruction: $I$
Landmarks:
which is shown in Fig. 1 (upper left). We constrain ChatGPT not to generate abstract landmarks and landmarks not mentioned in the instruction by stating it explicitly in the prompt.
Figure 3: Numbers of Landmarks and Cooccurrences.
Landmark Cooccurrence Generation. Based on the extracted navigational sequential landmarks, we can further obtain landmark cooccurrence priors from ChatGPT. Specifically, we pre-define 2 task examples for ChatGPT to generate cooccurrences. One task example followed by the prompt is as follows:
Tell me 3 co-occurrences of bedroom:
1.bed;
2.mirror;
3.nightstand;
Tell me 3 co-occurrences of $U_{k}^{la}$:
where $U_{k}^{la}$ is the $k$-th landmark in an instruction. We constrain ChatGPT to generate visible objects or scenes rather than abstract objects (e.g., wind) as cooccurrences by stating it explicitly in the prompt. We only show 3 cooccurrences due to the limited space. The whole prompts with all task examples for landmark extraction and landmark cooccurrence generation are given in the Supplementary Material.
Fig. 3 presents a visualization of number comparison between landmarks and cooccurrences. From Fig. 3 we can find that by introducing cooccurrences for landmark discovery, the agent can learn to align more diverse objects with observations and therefore capture rich alignment knowledge.
D.2 Correctable Landmark Discovery
Due to the lack of visual constraints, the obtained priors may be inconsistent with actual observations and cause misleading. To address this issue, we introduce a CLIP-driven correctable landmark discovery scheme to correct the priors in a learnable way according to actual observations, which contains the landmark shifting, the landmark discovery, and the learnable cooccurrence scoring. For the instruction $I$, let the landmark cooccurrence priors be $U=\{U^{la},U^{co}\}$, where $U^{la}=\{U^{la}_{k}\}_{k=1}^{N^{la}}$ and $U^{co}=\{\{U_{1,i}^{co}\}_{i=1}^{N^{co}},...,\{U_{N^{la},i}^{co}\}_{i=1}^{N^{co}}\}$ represent the extracted landmark list and cooccurrences, respectively. $N^{la}$ and $N^{co}$ are the numbers of landmarks and cooccurrences, respectively.
D.2.1 Landmark Shifting
At each timestep $t$, before the landmark discovery, we introduce a landmark shifting module based on CLIP to decide the current important landmark $U^{la}_{t}$ from the landmark list $U^{la}$. Since $U^{la}$ already implies the intended order to be searched, we keep a shifting pointer $z$ for extracting two adjacent landmarks $\{U_{z}^{la},U_{z+1}^{la}\}$ from $U^{la}$. The pointer $z=1$ when $t=0$. The landmark shifting module predicts the current important landmark $U^{la}_{t}$ from $\{U_{z}^{la},U_{z+1}^{la}\}$ based on the observations $O_{t}=\{O_{t,n}\}_{n=1}^{N_{o}}$. For each image $B_{t,n}$ in the single-view observation $O_{t,n}$, we first use the CLIP image encoder $E^{V}_{\mathrm{CLIP}}(\cdot)$ to obtain the feature $\mathbf{f}_{B_{t,n}}$. And we obtain the current observation feature $\mathbf{f}_{B_{t}}$ by an average operation:
$\mathbf{f}_{B_{t}}=\frac{1}{N_{o}}\sum_{n}\mathbf{f}_{B_{t,n}}=\frac{1}{N_{o}}\sum_{n}{E^{V}_{\mathrm{CLIP}}(B_{t,n})}.$ (7)
Then, we obtain the textual features $\{\mathbf{f}_{U_{z}^{la}},\mathbf{f}_{U_{z+1}^{la}}\}$ via the CLIP text encoder $E^{T}_{\mathrm{CLIP}}(\cdot)$:
$\mathbf{f}_{U_{z}^{la}}=E^{T}_{\mathrm{CLIP}}(U_{z}^{la}),\quad\mathbf{f}_{U_{z+1}^{la}}=E^{T}_{\mathrm{CLIP}}(U_{z+1}^{la}).$ (8)
With $\mathbf{f}_{B_{t}}$ and $\mathbf{f}_{U_{z}^{la}}$, we obtain the probability $p_{U_{z}^{la}}$ that the current important landmark $U^{la}_{t}$ is $U_{z}^{la}$ by:
$p_{U_{z}^{la}}=\frac{\mathrm{exp}(\mathrm{sim}(\mathbf{f}_{B_{t}},\mathbf{f}_{U_{z}^{la}})/\tau)}{\mathrm{exp}(\mathrm{sim}(\mathbf{f}_{B_{t}},\mathbf{f}_{U_{z}^{la}})/\tau)+\mathrm{exp}(\mathrm{sim}(\mathbf{f}_{B_{t}},\mathbf{f}_{U_{z+1}^{la}})/\tau)},$ (9)
where $\tau$ is the temperature parameter, $\mathrm{exp}(\cdot)$ represents the exponent operation, and $\mathrm{sim}(\cdot,\cdot)$ denotes the similarity computation, which is the dot product between two features. The probability $p_{U_{z+1}^{la}}$ is calculated similarly. Then $U^{la}_{t}=U_{z}^{la}$ if $p_{U_{z}^{la}}>p_{U_{z+1}^{la}}$ and $U^{la}_{t}=U_{z+1}^{la}$ if $p_{U_{z}^{la}}\leq p_{U_{z+1}^{la}}$. If $U^{la}_{t}=U_{z+1}^{la}$, the pointer $z$ will move to $z+1$ at timestep $t+1$. Due to the scene complexity, the pointer $z$ may not move during many navigation steps. Therefore, we force the movement of $z$ when it continuously points to the same position (see Fig. 4 for example). Specifically, we define a counter to record the number of no-shifting steps. If the number of no-shifting steps exceeds a pre-defined step threshold, we force the pointer $z$ to move to $z+1$ and reset the counter to 0. The step threshold is set to 1 when an instruction contains more than 3 landmarks. Otherwise, the step threshold is set as 2.
Note that Hong et al. also introduce a sub-instruction shifting module [48] which is learnable to facilitate the alignment between instructions and trajectories. Different from [48], we do not add learnable parameters in our heuristic landmark shifting module and instead construct a learnable cooccurrence scoring module (Sec. D.2.3) that effectively facilitates cooccurrence-based correctable landmark discovery. Moreover, we conduct fine-grained shifting among landmarks rather than sub-instructions to enable sequential landmark discovery.
D.2.2 Landmark Discovery
In discrete VLN task, the actions are chosen from candidate single-view observations. Therefore, to adapt our landmark discovery framework for navigation action prediction, we transform the landmark discovery as locating the landmark in one of the candidate observations, which is consistent with the action prediction process. With the current important landmark $U^{la}_{t}$ and its cooccurrences $\{U^{co}_{t,i}\}_{i=1}^{N^{co}}$, we conduct the landmark discovery through CLIP to highlight the single-view observation where $U^{la}_{t}$ with $\{U^{co}_{t,i}\}_{i=1}^{N^{co}}$ is most likely to appear for assisting action decision. To this end, for each single-view observation $O_{t,n}$, we calculate the probability $p_{U^{la}_{t,n}}$ that the current important landmark $U^{la}_{t}$ appears in it by:
$p_{U^{la}_{t,n}}=\frac{\mathrm{exp}(\mathrm{sim}(\mathbf{f}_{U^{la}_{t}},\mathbf{f}_{B_{t,n}})/\tau)}{\sum_{m=1}^{N_{o}}\mathrm{exp}(\mathrm{sim}(\mathbf{f}_{U^{la}_{t}},\mathbf{f}_{B_{t,m}})/\tau)},$ (10)
where $\mathbf{f}_{U_{t}^{la}}=\mathbf{f}_{U_{z}^{la}}$ or $\mathbf{f}_{U_{z+1}^{la}}$, $\mathbf{f}_{B_{t,n}}$ is the visual feature for $O_{t,n}$ (Eq. 7), $\mathrm{exp}(\cdot)$ represents the exponent operation, and $\mathrm{sim}(\cdot,\cdot)$ denotes the similarity computation, which is the dot product between two features. We also calculate the probability $p_{U^{co}_{t,i,n}}$ that $U^{co}_{t,i}$ appears in $O_{t,n}$ for each cooccurrence $U^{co}_{t,i}$ like Eq. 10.
D.2.3 Learnable Cooccurrence Scoring
Since the cooccurrence priors are extracted without visual constraints, the cooccurrences may not appear in the observations or may appear in multiple single-view observations to cause misleading. Therefore, we introduce a learnable cooccurrence scoring module to correct the priors according to actual observations. Moreover, due to the complexity of the scene and the instruction, it is hard for the large model to determine the correct landmark totally correctly. Through the introduction of the learnable cooccurrence scoring module trained together with the baseline agent, the noise brought by the landmark shifting module and the landmark discovery module can be effectively mitigated.
Specifically, we obtain the landmark score $s^{la}$ and the cooccurrence score $\{s_{i}^{co}\}_{i=1}^{N^{co}}$ through a dynamic updated state feature $\mathbf{f}_{S_{t}}$. Then we use $s^{la}$ and $s_{i}^{co}$ to obtain the corrected landmark prediction probability $\tilde{p}_{U_{t,n}}$ based on the probabilities $p_{U^{la}_{t,n}}$ and $\{\{p_{U^{co}_{t,1,n}}\}_{n=1}^{N_{o}},...,\{p_{U^{co}_{t,N^{co},n}}\}_{n=1}^{N_{o}}\}$. Through $s^{la}$ and $s_{i}^{co}$, the important cooccurrence in the priors can be highlighted while the unimportant ones are suppressed by $\tilde{p}_{U_{t,n}}$. At timestep $t$, we calculate $\mathbf{f}_{S_{t}}$ based on the instruction feature $\mathbf{f}_{I}$ (Eq. 1) and the current observation feature $\mathbf{f}_{B_{t}}$ (Eq. 7):
$\mathbf{f}_{S_{t}}=E^{l}([\mathbf{f}_{I}^{cls};\mathbf{f}_{B_{t}}]),$ (11)
where $\mathbf{f}_{I}^{cls}$ is the instruction feature of the [CLS] token in the baseline [16, 24], and $E^{l}(\cdot)$ contains a linear layer, a ReLU activation function, a layer normalization function, and a Dropout operation. Then, we obtain the cooccurrence score $s_{i}^{co}$ for cooccurrence $U^{co}_{t,i}$ ($1\leq i\leq N^{co}$) by:
$s_{i}^{co}=\mathrm{sim}(\mathbf{f}_{S_{t}},\mathbf{f}_{U^{co}_{t,i}}),$ (12)
where $\mathrm{sim}(\cdot,\cdot)$ denotes the similarity computation which is the dot product between two features and $\mathbf{f}_{U^{co}_{t,i}}=E^{T}_{\mathrm{CLIP}}(U^{co}_{t,i})$ is the textual feature. We also calculate the score $s^{la}$ for the current important landmark $U^{la}_{t}$ as Eq. 12. Based on the original probability $p_{U^{la}_{t,n}}$ and $\{p_{U^{co}_{t,i,n}}\}_{i=1}^{N^{co}}$ , we obtain the corrected landmark prediction probability $\tilde{p}_{U_{t,n}}$ for the single-view observation $O_{t,n}$ by:
$\tilde{p}_{U_{t,n}}=p_{U^{la}_{t,n}}\cdot s^{la}+\sum_{i=1}^{N^{co}}p_{U^{co}_{t,i,n}}\cdot s_{i}^{co},$ (13)
Consistency Loss. To constrain the optimization of the learnable cooccurrence scoring module, we introduce a consistency loss $\mathcal{L}_{\mathrm{cs}}$, which is the cross-entropy loss and calculated based on the corrected landmark prediction probability $\tilde{p}_{U_{t,n}}$ and the ground-truth action $g_{t}$:
$\mathcal{L}_{\mathrm{cs}}=-\sum_{t}g_{t}\mathrm{log}(\mathbf{\tilde{p}}_{U_{t}}),$ (14)
where $\mathbf{\tilde{p}}_{U_{t}}=\{\tilde{p}_{U_{t,n}}\}_{n=1}^{N_{o}}$. Since both action prediction and landmark discovery aim at selecting the ground-truth candidate observation, by enforcing the consistency between $\mathbf{\tilde{p}}_{U_{t}}$ and $g_{t}$, the scoring module can learn to correct the importance of cooccurrences for accurate landmark discovery. As shown in Fig. 2, through the learnable cooccurrence scoring module under the constraint of the consistency loss $\mathcal{L}_{\mathrm{cs}}$, the cooccurrence “counter” is significantly highlighted to indicate the ground-truth action.
D.3 Observation Enhancement
To encourage VLN agents to learn sufficient open-world alignment knowledge under the correctable landmark discovery framework, we design an observation enhancement strategy, where we use the corrected landmark feature $\mathbf{f}_{U_{t}}=\{\mathbf{f}_{U_{t,n}}\}_{n=1}^{N_{o}}$ and the original observation feature $\mathbf{f}_{O_{t}}=\{\mathbf{f}_{O_{t,n}}\}_{n=1}^{N_{o}}$ to obtain the enhanced observation features $\mathbf{f^{\prime}}_{O_{t}}=\{\mathbf{f^{\prime}}_{O_{t,n}}\}_{n=1}^{N_{o}}$ for action decision. Specifically, at timestep $t$, we calculate $\mathbf{f}_{U_{t,n}}$ by:
$\mathbf{f}_{U_{t,n}}=p_{U^{la}_{t,n}}\cdot s^{la}\cdot\mathbf{f}_{U^{la}_{t}}+\sum_{i=1}^{N^{co}}p_{U^{co}_{t,i,n}}\cdot s_{i}^{co}\cdot\mathbf{f}_{U^{co}_{t,i}},$ (15)
where $\mathbf{f}_{U^{la}_{t}}$ and $\mathbf{f}_{U^{co}_{t,i}}$ are the textual features of the landmark $U^{la}_{t}$ and the cooccurrence $U^{co}_{t,i}$. Through $s^{la}$ and $s_{i}^{co}$, the feature of the important cooccurrence can be highlighted. Then, we use $\mathbf{f}_{U_{t,n}}$ and $\mathbf{f}_{O_{t,n}}$ to obtain the enhanced observation feature $\mathbf{f^{\prime}}_{O_{t,n}}$ for $O_{t,n}$:
$\mathbf{f^{\prime}}_{O_{t,n}}=\mathbf{f}_{O_{t,n}}+\mathbf{f}_{U_{t,n}}.$ (16)
The enhanced observation features $\mathbf{f^{\prime}}_{O_{t}}=\{\mathbf{f^{\prime}}_{O_{t,n}}\}_{n=1}^{N_{o}}$ are finally utilized for action decision.
Contrastive Loss. For improving the modality alignment and encouraging the enhanced observation feature $\mathbf{f^{\prime}}_{O_{t,n}}$ to be more discriminative for facilitating action decision, we introduce a contrastive loss $\mathcal{L}_{\mathrm{ct}}$. Through $\mathcal{L}_{\mathrm{ct}}$, we enforce the observation feature $\mathbf{f}_{O_{t,n}}$ to be close with the paired corrected landmark feature $\mathbf{f}_{U_{t,n}}$ while far away from the non-paired corrected landmark feature $\overline{\mathbf{f}}_{U_{t,n}}=\{\mathbf{f}_{U_{t,j}}\}$ ($1\leq j\leq N_{o}$ and $j\neq n$). $\mathcal{L}_{\mathrm{ct}}$ is composed of two losses $\mathcal{L}_{\mathrm{ct}}^{o\rightarrow{u}}$ and $\mathcal{L}_{\mathrm{ct}}^{u\rightarrow{o}}$. We calculate $\mathcal{L}_{\mathrm{ct}}^{o\rightarrow{u}}$ by:
$\mathcal{L}_{\mathrm{ct}}^{o\rightarrow{u}}=-\sum_{t,n}\mathrm{log}(\frac{\mathrm{e}^{\mathrm{sim}(\mathbf{f}_{O_{t,n}},\mathbf{f}_{U_{t,n}})/\tau}}{\mathrm{e}^{\mathrm{sim}(\mathbf{f}_{O_{t,n}},\mathbf{f}_{U_{t,n}})/\tau}+\sum\limits_{\overline{\mathbf{f}}_{U_{t,n}}}\mathrm{e}^{\mathrm{sim}(\mathbf{f}_{O_{t,n}},\overline{\mathbf{f}}_{U_{t,n}})/\tau}}),$ (17)
where $\mathrm{sim}(\cdot,\cdot)$ denotes the similarity computation, which is the dot product between two features. $\mathcal{L}_{\mathrm{ct}}^{u\rightarrow{o}}$ is calculated in a symmetric way as $\mathcal{L}_{\mathrm{ct}}^{o\rightarrow{u}}$. Then we obtain the contrastive loss $\mathcal{L}_{\mathrm{ct}}$ by $\mathcal{L}_{\mathrm{ct}}=\frac{1}{2}(\mathcal{L}_{\mathrm{ct}}^{o\rightarrow{u}}+\mathcal{L}_{\mathrm{ct}}^{u\rightarrow{o}})$.
D.4 Action Prediction
The action $\mathbf{\tilde{a}}_{t}$ is predicted based on the enhanced observation feature $\mathbf{f^{\prime}}_{O_{t}}$ and the instruction feature $\mathbf{f}_{I}$ through the cross-modal Transformer encoder $E^{c}(\cdot)$ and the action prediction module $E^{a}(\cdot)$ (see Eq. 2 and Eq. 3). The total training objective $\mathcal{L}$ of CONSOLE is:
$\mathcal{L}=\mathcal{L}_{\mathrm{nav}}+\lambda_{1}\mathcal{L}_{\mathrm{cs}}+\lambda_{2}\mathcal{L}_{\mathrm{ct}},$ (18)
where $\lambda_{1}$ and $\lambda_{2}$ are the balance meta-parameters.
E Experiments
TABLE II: Comparison with SOTA methods on R2R. * denotes using additional large-scale unannotated data. Bold and Blue denote the best and runner-up results.
Method Val Seen Val Unseen Test Unseen
TL NE $\downarrow$ SR $\uparrow$ SPL $\uparrow$ TL NE $\downarrow$ SR $\uparrow$ SPL $\uparrow$ TL NE $\downarrow$ SR $\uparrow$ SPL $\uparrow$
Seq2Seq [1] 11.33 6.01 39 - 8.39 7.81 22 - 8.13 7.85 20 18
RCM+SIL(train) [26] 10.65 3.53 67 - 11.46 6.09 43 - 11.97 6.12 43 38
EnvDropout [6] 11.00 3.99 62 59 10.70 5.22 52 48 11.66 5.23 51 47
RelGraph [13] 10.13 3.47 67 65 9.99 4.73 57 53 10.29 4.75 55 52
PREVALENT [49] 10.32 3.67 69 65 10.19 4.71 58 53 10.51 5.30 54 51
ORIST [31] - - - - 10.90 4.72 57 51 11.31 5.10 57 52
[…]
HOP [33] 11.51 2.46 76 70 12.52 3.79 64 57 13.29 3.87 64 58
VLN-SIG [50] - - - - - - 72 62 - - 72 60
KERM [51] 12.16 2.19 80 74 13.54 3.22 72 61 14.60 3.61 70 59
Meta-Explore [52] 11.95 2.11 81 75 13.09 3.22 72 62 14.25 3.57 71 61
ScaleVLN* [53] 13.24 2.12 81 75 14.09 2.09 81 70 13.92 2.27 80 70
NavGPT [54] - - - - 11.45 6.46 34 29 - - - -
HAMT [16] (baseline) 11.15 2.51 76 72 11.46 3.62 66 61 12.27 3.93 65 60
HAMT+CONSOLE (ours) 11.73 2.65 74 70 11.87 3.39 69 63 12.42 3.78 66 61
DUET [24] (baseline) 12.32 2.28 79 73 13.94 3.31 72 60 14.73 3.65 69 59
DUET+CONSOLE (ours) 12.74 2.17 79 73 13.59 3.00 73 63 14.31 3.30 72 61
E.1 Experimental Setup
Datasets. We evaluate CONSOLE on four mainstream VLN benchmarks: R2R [1], REVERIE [2], R4R [3], and RxR [4]. R2R contains 90 indoor scenes with 7189 trajectories. REVERIE replaces the fine-grained instructions in R2R with high-level instructions to locate remote objects. R4R concatenates two adjacent tail-to-head trajectories in R2R, forming longer instructions and trajectories. RxR contains much more complex instructions and trajectories than R2R. Since CLIP [18] is pretrained on English language data, we use the English subset of RxR (both en-IN and en-US) for verification, which includes 26464, 2939, 4551 path-instruction pairs for Training, Val Seen, and Val Unseen, respectively.
Evaluation Metrics. The evaluation metrics for R2R [1] are: 1) Trajectory Length (TL): the average length of the agent’s navigated path in meters, 2) Navigation Error (NE): the average distance in meters between the agent’s destination and the target viewpoint, 3) Success Rate (SR): the ratio of success, where the agent stops within three meters of the target point, and 4) Success rate weighted by Path Length (SPL): success rate normalized by the ratio between the length of the shortest path and the predicted path. For R4R [3] and RxR [4], we further employ 5) the Coverage weighted by Length Score (CLS), 6) the normalized Dynamic Time Warping (nDTW), and 7) the Success weighted by nDTW (SDTW) for measuring the path fidelity. For REVERIE [2], three other metrics are used: 8) Remote Grounding Success Rate (RGS): the ratio of grounding the correct object when stopping, 9) Remote Grounding Success rate weighted by Path Length (RGSPL): weight RGS by TL, and 10) Oracle Success Rate (OSR): the ratio of containing a viewpoint along the path where the target object is visible.
Implementation Details. We conduct all experiments on an NVIDIA 3090 GPU. The batch size is set to 4 and the model is trained for 300K iterations on all datasets. The navigation loss of CONSOLE is the same as that in each baseline. The temperature parameter $\tau$ is set as 0.5 empirically. We use the SGD optimizer with the learning rate 0.1 for training $E^{l}(\cdot)$ in the learnable cooccurrence scoring module. The loss weights $\lambda_{1}$ and $\lambda_{2}$ are set as 0.1 on all datasets. The same augmented data in [16] is used for R2R for fair comparison. To sufficiently utilize the navigation ability of existing VLN agents, we train CONSOLE using the pretrained checkpoint of the baseline.
For facilitating implementation, we use the CLIP visual features (ViT-B/32) released by [16] instead of using the CLIP image encoder. And we pre-extract the landmark cooccurrence priors before training. To better activate the landmark discovery capability of CLIP, we add the prompt “a photo of a” like [18] to obtain the CLIP textual feature for a specific landmark during the landmark discovery. For the landmark extraction, we use the same task examples for R2R, R4R and RxR since the instructions for them are all fine-grained. For REVERIE which have high-level instructions, we pre-define different task examples. Since the number of landmarks in R2R augmented data is $\sim$75 times larger than that of R2R Train split, using ChatGPT directly for cooccurrence extraction has low feasibility regarding both time and cost. Alternatively, we obtain the cooccurrences for the augmented data through a pre-built landmark-cooccurrence vocabulary based on the landmark cooccurrence priors of R2R Train split. We compute the most semantically similar landmark in the vocabulary for each landmark in R2R augmentation data to ensure the quality of the extracted cooccurrences. For each path in R4R, we directly combine the landmarks and cooccurrences from 2 corresponding R2R instructions for improving efficiency.
TABLE III: Navigation and object grounding performance on REVERIE. Bold and underline denote the best and runner-up results.
Method Val Seen Val Unseen Test Unseen
[…]
Seq2Seq [1] 29.59 35.70 24.01 18.97 14.96 4.20 8.07 2.84 2.16 1.63 3.99 6.88 3.09 2.00 1.58
RCM [26] 23.33 29.44 21.82 16.23 15.36 9.29 14.23 6.97 4.89 3.89 7.84 11.68 6.67 3.67 3.14
SMNA [12] 41.25 43.29 39.61 30.07 28.98 8.15 11.28 6.44 4.54 3.61 5.80 8.39 4.53 3.10 2.39
FAST-MATTN [2] 50.53 55.17 45.50 31.97 29.66 14.40 28.20 7.19 7.84 4.67 19.88 30.63 11.60 11.28 6.08
SIA [55] 61.91 65.85 57.08 45.96 42.65 31.53 44.67 16.28 22.41 11.56 30.80 44.56 14.85 19.02 9.20
[…]
HOP [33] 54.81 56.08 48.05 40.55 35.79 30.39 35.30 25.10 18.23 15.31 29.12 32.26 23.37 17.13 13.90
TD-STP [32] - - - - - 34.88 39.48 27.32 21.16 16.56 35.89 40.26 27.51 19.88 15.40
KERM [51] 71.89 74.49 64.04 57.55 51.22 49.02 53.65 34.83 33.97 24.14 52.26 57.44 37.46 32.69 23.15
DUET [24] 71.75 73.86 63.94 57.41 51.14 46.98 51.07 33.73 32.15 23.03 52.51 56.91 36.06 31.88 22.06
CONSOLE (ours) 74.14 76.25 65.15 60.08 52.69 50.07 54.25 34.40 34.05 23.33 55.13 59.60 37.13 33.18 22.25
E.2 Quantitative Results
TABLE IV: Comparison on R4R val unseen split.
* denotes the results of our re-implementation.
[…]
SF [7] 8.47 24 30 - -
RCM [26] - 29 35 30 13
PTA [56] 8.25 24 37 32 10
EGP [14] 8.0 30.2 44.4 37.4 17.5
RelGraph [13] 7.43 36 41 47 34
[…]
HAMT [16] 6.09 44.6 57.7 50.3 31.8
HAMT* [16] 6.25 41.50 58.56 51.50 30.69
CONSOLE (ours) 5.99 43.22 61.45 54.34 32.72
TABLE V: Comparison on RxR-En val unseen split.
* denotes the results of our re-implementation.
[…]
EnvDrop [6] 38.5 34 54 51 32
Syntax [57] 39.2 35 56 52 32
HAMT* [16] 43.20 40.41 59.57 55.52 36.39
CONSOLE (ours) 46.34 42.92 60.85 57.11 38.85
Comparison with SOTAs. Table II [^1] The original value 2.29 of NE under Val Unseen in HAMT is a typo, which is actually 3.62 and confirmed by one of HAMT’s authors., Table III, Table IV, and Table V show the performance comparison between recent approaches and CONSOLE, where we can find that CONSOLE outperforms the strong baselines HAMT [16] and DUET [24] consistently on R2R, REVERIE, R4R, and RxR especially in unseen scenarios (e.g., CONSOLE outperforms DUET in SPL by 3% and 2% under Val Unseen and Test Unseen on R2R, respectively). The results demonstrate the strong generalization ability of CONSOLE for diverse instructions and scenarios. Moreover, CONSOLE establishes new SOTA results on R2R and R4R under unseen scenes. CONSOLE also shows good performance in seen scenarios on different datasets, e.g., it significantly outperforms DUET in Val Seen on REVERIE. All these results show the effectiveness of CONSOLE, demonstrating that CONSOLE successfully activates and utilizes open-world knowledge from large models for assisting VLN.
Note that although the concurrent work ScaleVLN [53] outperforms previous works as well as CONSOLE, it leverages significantly larger data resources and therefore a direct comparison would not be fair. Another concurrent work, NavGPT [54], introduces LLM to serve as the navigation decision backbone directly. Different from NavGPT [54], our CONSOLE combines the advantages of both LLM and existing vision-based VLN models to improve the navigation performance.
TABLE VI: Ablation study of the learnable cooccurrence scoring module (LS) and the observation enhancement module (OE) for different baselines (HAMT [16] & DUET [24]) and datasets (R2R [1] & REVERIE [2]).
Method R2R-HAMT R2R-DUET REVERIE-DUET
[…]
Baseline 3.62 66 61 3.31 72 60 46.98 33.73 51.07 32.15 23.03
No LS & No OE 3.53 67.39 62.42 3.12 72.54 60.64 48.68 33.27 54.96 32.83 22.53
LS & No OE 3.46 68.03 62.59 3.07 73.35 61.51 49.42 34.33 53.82 33.51 23.32
LS & OE 3.39 68.62 63.06 3.00 73.44 63.02 50.07 34.40 54.25 34.05 23.33
TABLE VII: Ablation Studies of cooccurrence number, landmark shifting module, feature fusion strategies, loss weights, and temperature parameter on R2R.
Method Val Unseen
[…]
Baseline [16] 3.62 66 61
$\triangleright$ Cooccurrence number:
num=0 3.60 66.84 61.78
num=1 3.60 66.45 61.62
num=5 3.53 67.39 62.42
num=10 3.52 67.31 62.19
$\triangleright$ Landmark shifting:
Global 3.59 66.88 61.51
Noforcedshift 3.54 67.22 62.05
Shift 3.53 67.39 62.42
$\triangleright$ strategies for fusing landmark features and observation features:
concatenation 3.61 66.24 61.12
summation 3.46 68.03 62.59
$\triangleright$ Loss weights:
$\lambda_{1}$=0.1 & $\lambda_{2}$=1 3.50 67.48 62.56
$\lambda_{1}$=1 & $\lambda_{2}$=0.1 3.63 67.35 62.53
$\lambda_{1}$=1 & $\lambda_{2}$=1 3.58 67.65 62.26
$\lambda_{1}$=0.1 & $\lambda_{2}$=0.1 (ours) 3.39 68.62 63.06
$\triangleright$ Temperature parameter $\tau$:
$\tau$=0.05 3.48 67.56 62.36
$\tau$=0.1 3.44 68.03 62.71
$\tau$=1 3.50 68.11 63.10
$\tau$=0.5 (ours) 3.39 68.62 63.06
TABLE VIII: Comparison of LLMs on R2R. The baseline we choose is DUET [24].
Method Val Unseen
[…]
Baseline [16] 3.31 72 60
Spacy 3.11 73.05 61.33
GPT-3 3.00 73.82 61.56
Vicuna7B 3.06 73.31 61.65
Vicuna13B 3.07 73.39 62.14
ChatGPT (ours) 3.00 73.44 63.02
Ablation Study. To further analyze the effects of different components in CONSOLE, we conduct extensive ablation studies, and the results are given in Table VI and Table VII. Specifically, to analyze the effects of two core modules in CONSOLE, i.e., the learnable cooccurrence scoring module (LS) and the observation enhancement module (OE), we give the ablation study results of them for different baselines and datasets in Table VI. Concretely, “LS & No OE” outperforms “No LS & No OE” by generally  0.8% improvement on different baselines and datasets, which shows the effectiveness of learnable cooccurrence scoring module. For example, “LS & No OE” outperforms “No LS & No OE” by 1.06% and 0.87% on SPL for REVERIE-DUET and R2R-DUET, respectively. Similarly, “LS & OE” outperforms “LS & No OE” on all metrics on different baselines and datasets, which also shows the effectiveness and the generalization ability of the observation enhancement module. The results in Table VI sufficiently show the effectiveness of two core modules in CONSOLE.
We further conduct the ablation studies for the cooccurrence number, landmark shifting module, strategies for fusing the corrected landmark features and the original observation features, loss weights, and temperature parameter, and the results are presented in Table VII. From Table VII we can draw into the following conclusions: 1) The results of “Cooccurrence number” show that the landmark cooccurrence priors greatly facilitate the landmark discovery, e.g., “num=5” outperforms “num=0” by 0.55% and 0.64% in SR and SPL, respectively. Moreover, the results of “num=5,10” show that adding too many co-occurrences may not bring further performance improvement. It is reasonable since too many co-occurrences may introduce unexpected noise to mislead the agent in the meanwhile. 2) The results of “Landmark Shifting” show that landmark shifting module is helpful for accurate landmark discovery. Especially, “Shift” outperforms “Global” by 0.91% in SPL. 3) The results of “strategies for fusing the corrected landmark features and the original observation features” show that “summation” outperforms “concatenation” and “concatenation” is comparable with the baseline. This result reveals that “summation” is a more effective way for fusing the corrected landmark features and the original observation features. 4) The results of “Loss weights” show that all settings with different scales of the loss weights $\lambda_{1}$ and $\lambda_{2}$ outperform the baseline, demonstrating that our method is insensitive to the loss weight. Moreover, our current setting ($\lambda_{1}=0.1$ and $\lambda_{2}=0.1$) achieves the best performance among other loss weight settings, demonstrating that selecting the proper loss weight to balance the scale of different loss items is helpful for improving the performance. 5) The results of the temperature parameter $\tau$ show that our method is not sensitive to the selection of the temperature parameter and the performance is best in most metrics when the temperature parameter is set to 0.5.
Comparison of different LLMs. We further conduct quantitative experiments to analyze how different LLMs impact the navigation performance. Specifically, we use different LLMs for navigational sequential landmark extraction, including ChatGPT (used by our CONSOLE), GPT-3 [17] (the predecessor of ChatGPT), Vicuna [58] (a recent strong open source alternative to ChatGPT which has two sizes——7B and 13B), and Spacy [59] (a NLP library), and feed the extracted landmarks into the consequent modules. We access the APIs of ChatGPT and Vicuna [58] in May, 2023. We access the API of GPT-3 [17] in Dec, 2022. The results are given in Table VIII. From Table VIII we can observe that: 1) ChatGPT outperforms other competitors in most metrics; 2) GPT-3, Vicuna13B, and Vicuna7B are comparable to ChatGPT in some metrics; 3) Vicuna13B outperforms Vicuna7B especially in SPL; 4) Spacy performs the worst among all competitors. These results show that more powerful LLMs can provide more accurate knowledge for navigation.
Figure 4: Visualization examples of action decision. The ground-truth (GT) action and the corrected landmark prediction are denoted in the green boxes.
Figure 5: Visualization example of landmark shifting. The ground-truth action (GT) and the mentioned landmark are annotated in the green boxes and orange boxes, respectively.
Figure 6: Visualization of comparison of different task examples and prompts.
Figure 7: Comparison of navigational sequential landmark extraction among different LLMs. The correct landmarks are denoted in orange. For extraction results, noise is marked in red, landmarks in the wrong order are marked in blue, and missing landmarks are marked in purple with brackets.
Figure 8: Comparison results of obtaining cooccurrences between ChatGPT and ConceptNet.
E.3 Visualization
In this subsection, we present rich visualization results to deeply analyze the effects of different modules in CONSOLE, including the learnable cooccurrence scoring module, the landmark shifting module, the customized designs of prompts, etc.
Action Decision. Fig. 4 gives some visualization examples of action decision, where we can find that CONSOLE effectively corrects the landmark cooccurrence priors for facilitating action prediction. For example, in Fig. 4(a), the importance weight of “carpet” which appears in “candidate_1” but does not appear in “candidate_5 (GT)” is significantly low. And the importance weights of “landing”, “steps”, and “handrail’ which appear in “candidate_5 (GT)” are relatively high. As a result, the landmark prediction is successfully corrected to indicate the GT action.
Landark shifting. Fig. 5 shows that our landmark shifting module takes better advantage of the landmark navigation order for landmark discovery. It can effectively mitigate the stuck and disordered landmark prediction caused by scene complexity. For example, in Fig. 5(a) since the mirror appears in the observations during multiple steps, “Noforcedshift” continuously predicts “mirror” as the current important landmark. CONSOLE, however, predicts correct landmarks at different timesteps.
Impact of task examples and prompts. Fig. 6 gives the visualization of landmark extraction and cooccurrence generation regarding different task examples and prompts. From Fig. 6 we can find that by providing task examples regarding diverse landmark descriptions, we can successfully improve the landmark extraction for LLM. For example, in Fig. 6(a), when prompted with only one example that extracts landmarks of a single noun, ChatGPT only extracts table. After adding the two examples that extract noun-phrase landmarks, table with china can be successfully extracted.
Moreover, we can also observe that adding the constraint in the prompt can effectively enable LLM to generate desired occurrences. For example, in Fig. 6(b), ChatGPT generates abstract words such as support without the constraint. However, through adding the constraint, ChatGPT generates more reliable cooccurrences like statue.
Navigational sequential landmark extraction. To verify the capability of existing LLMs for accurate landmark extraction, we give the visualization examples of landmark extraction comparison among different methods in Fig. 7. From Fig. 7, we can find that: (1) ChatGPT performs the best generally among all compared models; (2) Vicuna13B [58] is comparable to ChatGPT while sometimes extracting wrong navigational order; (3) Vicuna7B [58] also suffers from wrong navigational order (e.g., in Fig. 7(a), Vicuna7B extracts the “stone tiles” before extracting the “wodden floor”); (4) GPT-3 [17] is slightly inferior to ChatGPT for fine-grained landmark extraction (e.g., in Fig. 7(a), GPT-3 extracts the “wall” and the “fireplace” rather than extracting the “wall containing fireplace”); (5) Spacy [59] can only extract the textual order rather than the navigational order and extracts a lot of noise (e.g., in Fig. 7(b), Spacy extracts the “right”, the “rest”, and the “base”). These landmark extraction visualization results reveal the great potential of utilizing LLMs for assisting VLN. Moreover, it is also important to resort to more low-cost LLMs such as Vicuna [58] for helping embodied agents in the future.
Landmark cooccurrence extraction. To verify the superiority of LLM in landmark cooccurrence extraction, we present the comparison of obtaining landmark cooccurrences between ChatGPT and ConceptNet in Fig. 8. To get the cooccurrences of a target landmark through ConceptNet, we respectively request its related terms and nodes which are connected with relations of ’AtLocation’ and ’LocatedNear’. From Fig. 8 we can observe that LLM can generate visible cooccurrences as constrained in the prompt (Sec. D.1) which are informative and high-quality. ConceptNet, however, tends to generate synonyms and concepts rather than visible cooccurrences. For example, in Fig. 8 (a), for the target hall, ChatGPT generates multiple cooccurrences of hall like artwork and chandelier to help recognize the location of the hall. However, ConceptNet retrieves hallway and building which are uninformative.
Moreover, LLM can adapt to flexible inputs and process them robustly, whereas ConceptNet may generate poor results without preprocessing. In Fig. 8 (b), for the target base of stairs (landmark formed as a phrase), ChatGPT generates valid cooccurrences of stairs such as handrail. However, ConceptNet is unable to focus on the important part (stairs) and instead searches invalid related terms of base.
F Conclusion
In this paper, we propose CONSOLE for VLN, which casts VLN as an open-world sequential landmark discovery problem by introducing a correctable landmark discovery framework based on two powerful large models. We harvest rich landmark cooccurrence commonsense from ChatGPT and employ CLIP for conducting landmark discovery. A learnable cooccurrence scoring module is constructed to correct the priors provided by ChatGPT according to actual observations. Experimental results show that CONSOLE outperforms strong baselines consistently on R2R, REVERIE, R4R, and RxR. CONSOLE establishes new SOTA results on R2R and R4R under unseen scenarios.
We believe that our work can provide a meaningful reference to the researchers in both the VLN and the embodied AI areas in how to effectively harvest and utilize the helpful knowledge inside large models for assisting robotic tasks, which is a very promising way to improve the performance of robots. In future work, we would like to resort to more low-cost large models for assisting embodied AI tasks, which is more practical in real-world applications. It is also worth studying to develop efficient in-domain pretraining paradigms to adapt large models to embodied AI tasks.
Acknowledgments
This work was supported in part by National Science and Technology Major Project (2020AAA0109704), Guangdong Outstanding Youth Fund (Grant No. 2021B1515020061), Mobility Grant Award under Grant No. M-0461, Shenzhen Science and Technology Program (Grant No. GJHZ20220913142600001), Nansha Key RD Program under Grant No.2022ZD014, CAAI-Huawei MindSpore Open Fund. We thank MindSpore for the partial support of this work, which is a new deep learning computing framwork [^2] https://www.mindspore.cn/.